A charge -2c moved from point A to point X in an electric field.The work done is 6J and kinetic energy of charge reduces by 10J during the process.The potential difference Vₐ-Vₓ is?
OPTIONS:
8V
-8V
4V
-4V

Grade:12th Pass

1 Answers

Moinuddin
5 years ago
-8v 
Change in the every of the charge dusting the process=-6j-10j=-16j
Change in every = charge*(vx-va)
-16=-2(vx-va)
Vx-va=8
Va-vx=-8v